The equation for a distance, s(m), travelled in time t(s) by an object starting with an initial velocity u(ms-1) and uniform acceleration a(ms-2) is:




1
Expert's answer
2021-11-16T10:24:30-0500

Given distance traveled= s(m)

Time taken= t(s)

Initial velocity= u(ms-1)

Uniform acceleration= a(me-2)


The equation for distance traveled is given by the newton law that's states

"s=ut+\frac{1}{2}at^2"